1. Fault analysis in diagnostic assessment is a powerful way to establish whether a teacher understands the development of number sense and number concepts. Go through the following (true) case study of a Grade 7 learner with an extra mathematics teacher who tries to make sense of his ever increasing backlog:
“John, what is three fifths plus three fifths?” 3/5 +3/5
“I think it is six tenths Mam.”
“John, show me on this number line where three fifths is”
|______|______|______|______|______|______|______|__
0 1 2 3 4 5 * 6 7
“It is somewhere over here I think Mam.”

Analyse the learner’s fault(s) and try to establish which concepts are not in place.
1
Expert's answer
2020-05-07T15:55:16-0400

"\frac{3}{5} + \frac{3}{5}=\frac{6}{5}" , not "\frac{6}{10}" as John thinks.

On this number line three fifths is